# Mitigate Language Priors in Large Vision-Language Models by Cross-Images Contrastive Decoding

### v1

@@ IV-B Identifying Detrimental Priors
We employ Jensen-Shannon (JS) divergence to measure the consistency between two logit distributions produced by different images:
@@
where $p_{\theta}(y_{t}\mid v,x,y_{<t})$ is defined in Eq.1. The statistical results are visualized in Fig.5. We observe that the JSD values can be clearly divided into two distinct ranges: $\log_{10}(\mathcal{D}{\mathrm{JS}})>-4$ and $\log_{10}(\mathcal{D}_{\mathrm{JS}})\leq-4$. Larger JSD values correspond to detrimental priors, while smaller values indicate essential priors. Based on this observation, we set a threshold of $\gamma=-4$ to distinguish between the two types of priors.
IV-C Cross-Images Contrastive Decoding
The model performs regular decoding when essential priors are detected, and applies contrastive decoding to eliminate detrimental priors. For detrimental priors, we hypothesize that smaller distributional differences indicate stronger detrimental effects. Consequently, we define the contrastive coefficient $\alpha$ in Eq. 3 as:
$\alpha=1-\log_{10}(\mathcal{D}_{\mathrm{JS}})$ (6)
Based on the observed density of $\mathcal{D}_{\mathrm{JS}}$ values, we clip $\alpha$ to the interval $[1,3]$ to ensure numerical stability.
The final logit distribution is defined as follows:
$\mathrm{logit}_{CICD}(y_{t}\mid v,v^{\prime},x,y_{<t})=$ (7)
$\begin{cases}\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})&\text{if }\mathcal{D}_{\mathrm{JS}}\leq\gamma\\ \mathrm{logit}_{CD}(y_{t}\mid v,v^{\prime},x,y_{<t})&\text{if }\mathcal{D}_{\mathrm{JS}}>\gamma\end{cases}$
where $\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})$ is defined in Eq. 1, and $\mathrm{logit}_{CD}(y_{t}\mid v,v^{\prime},x,y_{<t})$ is defined in Eq. 3.
V Experiment
V-A Benchmarks
To evaluate the effectiveness of our method in eliminating language priors, we employ four multimodal hallucination benchmarks: two generative benchmarks (CHAIR [10] and DetailCaps [11]), one discriminative benchmark (POPE [14]), and one hybrid benchmark (AMBER [30]).
CHAIR assesses the proportion of hallucinated objects, which are generated by the model but not present in the reference annotations. Following prior works, we randomly select 500 images from the MSCOCO [46] dataset as the test set. We additionally select another 500 images to construct a validation set for hyperparameter tuning.
DetailCaps is a fine-grained image captioning benchmark, accompanied by ground-truth detail captions generated by GPT-4V [47], Gemini-1.5-Pro [48], and GPT-4o [49] for evaluation. It contains 4,870 images from multiple datasets, and we use a subset of 700 images from MSCOCO to conduct our experiments.
POPE is a widely adopted benchmark for evaluating object hallucinations by prompting LVLMs to identify whether a specific object is present in the image. It consists of 1,500 images, including 1,000 from MSCOCO and 500 from GQA [50].
AMBER combines generative and discriminative tasks, and is evaluated on a curated set of 1,004 images. In addition to image captioning, it includes 14,216 questions designed to assess hallucinations in object, attribute, and relation recognition.
V-B Evaluated LVLMs
We evaluate our method on six widely-used LVLMs: InstructBLIP [6], LLaVA-1.5 [31], LLaVA-Next [32], Qwen-VL-Chat [7], Qwen2-VL [33], and Qwen2.5-VL [8]. InstructBLIP and LLaVA-1.5 are based on Vicuna-7B [3] as their language backbone, while LLaVA-Next is built upon LLaMA3. The Qwen-VL series models, including Qwen2-VL and Qwen2.5-VL, are developed based on the QwenLM framework. Regarding vision encoders, InstructBLIP, LLaVA-1.5, LLaVA-Next, and Qwen-VL-Chat all use the pretrained CLIP model [29], whereas Qwen2-VL and Qwen2.5-VL employ a fine-tuned ViT [35]. All models are implemented at the 8B parameter scale.
TABLE I: Results on POPE. Acc. and F1 represent the average accuracy and F1-score across three datasets (MSCOCO, A-OKVQA, and GQA), respectively. The bolded results denote the best in each setting and the underlined results denote the second-best.
@@ V-B Evaluated LVLMs
V-C Baselines
We compare our method with several training-free baselines that specifically target language priors. VCD [21] introduces noise into the input image to suppress language priors, and then applies contrastive decoding for their elimination. ICD [22] constructs negative logits using adversarial instructions and enhances positive signals through contrastive decoding. IBD [23] strengthens the model’s attention to visual features, thereby increasing the influence of image information during decoding. DeFG [26] converts the generated caption into an image and re-generates the caption from that image, leveraging the preserved visual content to reduce hallucinations.
V-D Implementation Details
We use sampling-based decoding for next-token prediction with default settings. Following [21], we incorporate adaptive plausibility constraints into the decoding process. We use the MSCOCO 2014 validation set as the image retrieval corpus, which contains 40,504 images. For retrieval, we adopt CLIP111openai/clip-vit-large-patch14-336 as the feature extractor and perform similarity-based retrieval using precomputed image embeddings. All experiments are conducted on a single 24GB NVIDIA RTX 3090 GPU.
V-E Main Results
Results on POPE
POPE is a discriminative benchmark that incorporates three distinct datasets: MSCOCO, A-OKVQA, and GQA. Each dataset uses three different negative sampling settings: Random, Popular, and Adversarial. We conducted a comprehensive comparison between our method and other baselines across three distinct LVLMs, as detailed in Tab. I. For streamlined presentation, we report the average performance across all three datasets as the final result.
Our method demonstrates superior overall performance, particularly in the most critical metric, achieving the highest F1 scores across all experimental settings. This suggests that we can utilize cross-image consistency to effectively alleviate language priors and promote LVLMs to convey faithful visual information. Additionally, our method demonstrates consistently superior performance across diverse LVLMs, evidencing remarkable generalization capability. As a training-free approach, it offers immediate deployability in practical applications. Another key finding reveals that our method achieves significantly enhanced performance in the Popular setting, demonstrating an average F1 score improvement of 2.84 over the second-best approach. The Popular setting uses the most frequent objects to construct negative samples, which makes LVLMs more susceptible to language priors. The most substantial performance improvements achieved by our method under this setting further demonstrate its exceptional efficacy in mitigating language priors.
TABLE II: Results on the test set of CHAIR. MaxLen=512 indicates setting the maximum generation length of 512 tokens for the model. The bolded results denote the best in each setting and the underlined results denote the second-best.
Setting Method LLaVA-1.5 InstructBLIP
@@ V-E Main Results
MaxLen=512 Regular 54.6 16.4 72.6 106.8 62.6 19.5 66.9 113.6
@@
MaxLen=64 Regular 24.4 8.9 56.6 53.7 35.6 13.2 56.4 54.8
VCD 25.0 8.3 59 54.0 32.2 10.3 60.6 55.2
ICD 23.2 8.1 58.4 54.4 29.8 9.8 60.6 55.7
IBD 21.2 6.9 58.8 54.5 27.8 9.2 60.3 56.1
DeFG 22.4 7.2 58.2 54.4 32.4 11 59.4 56.0
CICD 18.0 6.1 59.6 53.3 23.8 7.7 62.2 54.8
TABLE III: Results on AMBER. We set the maximun generation length to 512 for the models in this benchmark. AMBER metirc is calculated by $(100-CHAIR+F1)/2$. The bolded results denote the best in each setting and the underlined results denote the second-best.
@@
Results on CHAIR
CHAIR is specifically designed to assess object hallucination in the image captioning task, which is one of the most language-prior-susceptible tasks in multimodal research. We evaluate our method in two settings with different maximum generation lengths; the results are shown in Tab. II.
Our method demonstrates superior performance across all experimental conditions, with particularly pronounced improvements observed at longer generation lengths. Longer context lengths amplify the model’s susceptibility to language priors, making these experimental results particularly compelling evidence of our method’s efficacy in mitigating such priors. Compared to baselines, our method’s principal advantage lies in its complete preservation of visual information, which is directly reflected in the superior experimental results. Additionally, we preserve the essential language priors, which are also important in generative tasks.
Results on AMBER
AMBER comprises generative tasks (image captioning) and discriminative tasks (image understanding). The results are shown in Tab. III. We can observe that our method achieves the lowest hallucination frequency in generative tasks while simultaneously attaining state-of-the-art performance on the comprehensive AMBER metric. This superior performance stems from our focused approach to language prior mitigation. This problem is particularly critical for generative tasks, where language priors manifest more severely than in discriminative tasks, explaining our method’s pronounced advantages in image captioning.
TABLE IV: Results on DetailCaps. The bolded results denote the best in each setting and the underlined results denote the second-best.
@@
LLaVA-1.5 InstructBLIP Qwen-VL-Chat
Regular 52.60 52.99 30.73
VCD 52.91 53.20 31.28
ICD 52.82 53.24 29.70
IBD 52.48 54.14 29.31
DeFG 52.72 53.06 30.21
CICD 55.80 54.20 31.34
Results on DetailCaps
We further validate our method’s effectiveness on the image captioning task, which is more susceptible to language priors. Tab. IV presents the evaluation results on DetailCaps. The CAPTURE metric quantitatively evaluates caption quality through both lexical matching and semantic alignment, measuring the consistency of key content (entities, attributes, and relations) between generated captions and multiple ground-truth references. Our method achieves the best performance across all three evaluated LVLMs, demonstrating its robust capability to mitigate language priors and significantly enhance image captioning performance.
TABLE V: Results on more LVLMs. F1 is the average F1-score across three datasets (MSCOCO, A-OKVQA, and GQA). The bolded results denote the best in each setting.
@@
LLaVA-next Regular 39.0 10.8 62.1 85.47 83.14 79.87
@@
Qwen2-VL Regular 53.8 10.0 69.4 88.33 85.82 82.43
CICD 35.0 9.2 60.5 90.67 88.07 84.53
Qwen2.5-VL Regular 41.2 11.5 65.7 87.83 85.66 83.12
CICD 37.2 9.3 60.0 89.95 88.00 84.75
V-F Effectiveness on More LVLMs
@@
To comprehensively demonstrate the effectiveness and generalizability of our method, we select three of the most recent open-source LVLMs to conduct experiments. The results are presented in Tab. V. Our method demonstrates significant performance improvements in all experiments, consistently reducing hallucinations across both generative and discriminative tasks through effective mitigation of language priors. The improvements observed across different LVLMs confirm our method’s practical advantages: an elegantly simple framework coupled with state-of-the-art effectiveness.
TABLE VI: This experiment conducted by LLaVA-1.5. Random Image and Retrieved Image represent the contrastive images selected through random sampling and image retrieval, respectively. w/o Image represents the contrastive distribution is constrcuted by removing the input image.

### v2

@@ 4.2 Identifying Detrimental Priors
We employ the Jensen–Shannon (JS) divergence to measure the consistency between two logit distributions produced from different images:
@@
where $p_{\theta}(y_{t}\mid v,x,y_{<t})$ is defined in Eq.1. The statistical results are visualized in Fig.5. We observe that the JSD values can be clearly separated into two distinct ranges: $\log_{10}(\mathcal{D}{\mathrm{JS}})>-4$ and $\log_{10}(\mathcal{D}_{\mathrm{JS}})\leq-4$. Higher JSD values are associated with detrimental priors, while lower values correspond to essential priors, as shown in Fig.2 and Fig.3. Based on this observation, we set a threshold of $\gamma=-4$ to distinguish between the two types of priors.
@@
Table 1: Results on POPE. Acc. and F1 represent the average accuracy and F1-score across three datasets (MSCOCO, A-OKVQA, and GQA), respectively. The bolded results denote the best in each setting, and the underlined results denote the second-best.
4.3 Cross-Images Contrastive Decoding
After completing the above preparation, we integrate the proposed method into the contrastive decoding process. Specifically, the model performs regular decoding when essential priors are detected and applies contrastive decoding to suppress detrimental priors. In the case of detrimental priors, smaller distributional differences imply a stronger detrimental effect. Accordingly, we define the contrastive coefficient $\alpha$ in Eq. 3 as: $\alpha=1-\log_{10}(\mathcal{D}_{\mathrm{JS}})$. Based on the observed density of $\mathcal{D}_{\mathrm{JS}}$ values, we clip $\alpha$ to the interval $[1,3]$ to ensure numerical stability.
The final logit distribution is defined as follows:
$\mathrm{logit}_{CICD}(y_{t}\mid v,x,y_{<t})=$ (6)
$\begin{cases}\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})&\text{if }\mathcal{D}_{\mathrm{JS}}\leq\gamma\\ \mathrm{logit}_{CD}(y_{t}\mid v,x,y_{<t})&\text{if }\mathcal{D}_{\mathrm{JS}}>\gamma\end{cases}$
where $\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})$ is defined in Eq. 1, and $\mathrm{logit}_{CD}(y_{t}\mid v,x,y_{<t})$ is defined in Eq. 3.
5 Experiment
MaxLen Method LLaVA-1.5 InstructBLIP
@@ 5 Experiment
64 Regular 24.4 8.9 56.6 53.7 35.6 13.2 56.4 54.8
VCD 25.0 8.3 59 54.0 32.2 10.3 60.6 55.2
ICD 23.2 8.1 58.4 54.4 29.8 9.8 60.6 55.7
IBD 21.2 6.9 58.8 54.5 27.8 9.2 60.3 56.1
DeFG 22.4 7.2 58.2 54.4 32.4 11 59.4 56.0
CICD 18.0 6.1 59.6 53.3 23.8 7.7 62.2 54.8
512 Regular 54.6 16.4 72.6 106.8 62.6 19.5 66.9 113.6
@@
Table 2: Results on the test set of CHAIR. MaxLen denotes the maximum generation length.
5.1 Experiment Settings
Due to space constraints, we present only the key aspects of our experimental setup here. Detailed settings can be found in Sec.B.
Benchmarks
We evaluate the performance of CICD across four widely adopted multimodal hallucination benchmarks to validate its effectiveness on different tasks. These include two generative benchmarks (CHAIR Rohrbach et al. (2018) and DetailCaps Ye et al. (2025)), one discriminative benchmark (POPE Li et al. (2023b)), and one hybrid benchmark (AMBER Wang et al. (2024a)).
Evaluated LVLMs
We evaluate the generalizability of CICD across six LVLMs from three model families: InstructBLIP Dai et al. (2023), the LLaVA series (LLaVA-1.5 Liu et al. (2024c) and LLaVA-Next Liu et al. (2024d)), and the Qwen series (Qwen-VL-Chat Bai et al. (2023b), Qwen2-VL-Instruct Wang et al. (2024b), and Qwen2.5-VL-Instruct Bai et al. (2025)). Unless stated otherwise, all models are implemented at the 7B scale.
Baselines
We compare CICD with four recent training-free methods that specifically target language priors: VCD Leng et al. (2024), ICD Wang et al. (2024c), IBD Zhu et al. (2024b), and DeFG Zhang et al. (2025). All of these methods construct negative visual contexts based on the original image. In contrast, we identify several limitations of this approach and propose CICD as a more effective alternative.
5.2 Implementation Details
We use sampling decoding for next-token prediction with default settings. Following Leng et al. (2024), we incorporate adaptive plausibility constraints into the decoding process. We use the MSCOCO 2014 validation set as the image retrieval corpus, which contains 40,504 images. All experiments are conducted on an NVIDIA RTX 3090 GPU.
@@ 5.2 Implementation Details
Table 3: Results on AMBER. The maximum generation length is 512. AMBER metirc is calculated by $(100-CHAIR+F1)/2$.
5.3 Main Results
Method LLaVA-1.5 InstructBLIP Qwen-VL-Chat
@@ 5.3 Main Results
Regular 55.7 17.4 52.60 58.6 18 52.99 9.9 7.3 30.73
VCD 55.7 16.8 52.91 59.6 18.9 53.20 7.3 5.9 31.28
ICD 53.9 16.6 52.82 55.7 16.7 53.24 5.6 7 29.70
IBD 54.1 15.8 52.48 56.4 15.5 54.14 4.6 2.9 29.31
DeFG 55.7 16.6 52.72 59.1 17.4 53.06 5.6 3.9 30.21
CICD 45.6 13.1 55.80 45.7 13.2 54.20 5.9 3.5 31.34
Table 4: Results on DetailCaps. The maximum generation length is 512.
Results on POPE
As shown in Tab. 1, our method achieves the highest F1 scores across all experimental settings. This suggests that we can utilize cross-image consistency to effectively alleviate language priors and promote LVLMs to convey faithful visual information. The Popular setting uses frequent objects for negative samples, increasing susceptibility to language priors. The strong performance gains of our method in this setting highlight its effectiveness in mitigating such priors.
@@
LLaVA-Next Regular 41.8 11.9 60.2 85.47 83.14 79.87
@@
Qwen2-VL-Instruct Regular 53.8 10.0 69.4 88.33 85.82 82.43
CICD 50.6 9.4 73.6 90.67 88.07 84.53
Qwen2.5-VL-Instruct Regular 41.2 11.5 65.7 87.83 85.66 83.12
CICD 39.4 9.6 68.1 89.95 88.00 84.75
Table 5: Results on more LVLMs. F1 is the average F1-score across three datasets (MSCOCO, A-OKVQA, and GQA). The bolded results denote the best in each setting.
Results on CHAIR
The results in Tab. 2 demonstrate our method’s superior performance across all experimental conditions, with particularly pronounced improvements observed at longer generation lengths. Longer context lengths amplify the model’s susceptibility to language priors, making these experimental results particularly compelling evidence of our method’s efficacy in mitigating such priors. Compared to baselines, our method’s principal advantage lies in its complete preservation of visual information and essential language priors, which is directly reflected in the superior performance in image captioning.
Results on AMBER
As illustrated in Tab. 3, our method achieves the lowest hallucination frequency in generative tasks while simultaneously attaining state-of-the-art performance on the comprehensive AMBER metric. Language prior is particularly critical for generative tasks, where language priors manifest more severely than in discriminative tasks, explaining our method’s pronounced advantages in image captioning.
Results on DetailCaps
Tab. 4 presents the evaluation results on DetailCaps. The CAPTURE metric quantitatively evaluates caption quality through both lexical matching and semantic alignment, measuring the consistency of key content (entities, attributes, and relations) between generated captions and multiple ground-truth references. Our method achieves the best performance across all three evaluated LVLMs, demonstrating its robust capability to mitigate language priors and significantly enhance image captioning performance.
5.4 Effectiveness on More LVLMs
@@
To comprehensively demonstrate the effectiveness and generalizability of our method, we select three of the most recent open-source LVLMs to conduct experiments. The results are presented in Tab. 5. Our method demonstrates significant performance improvements in all experiments, consistently reducing hallucinations across both generative and discriminative tasks through effective mitigation of language priors. We also test our method on these LVLMs by DetailCaps and the results are shown in Tab. 7. The improvements observed across different LVLMs confirm our method’s practical advantages: an elegantly simple framework coupled with state-of-the-art effectiveness.
Further experiments will be discussed in Appendix B.
6 Conclusion
In this paper, we address the issue of language priors in LVLMs. Through a comprehensive investigation, we reveal that language priors are intrinsic to the model and remain consistent across different images. Building on this observation, we propose Cross-Image Contrastive Decoding (CICD) — a training-free method that effectively mitigates language priors while preserving visual information. Extensive experiments across multiple benchmarks and LVLM architectures demonstrate the effectiveness of CICD, particularly in image captioning tasks where language priors are most prominent.
The simplicity and efficiency of CICD offer strong practical potential, enabling seamless integration with other hallucination mitigation methods to further enhance LVLM performance. This integration forms a key direction for our future work.
7 Limitation
We observe that language priors remain consistent across different images. Based on this insight, we propose the Cross-Images Contrastive Decoding (CICD) method to eliminate language priors. Our method has two main limitations compared to regular decoding. First, it requires an additional image that differs sufficiently from the original input. Second, both the original and the additional images must undergo separate forward passes to obtain their respective logit distributions for contrastive decoding.
Addtional Image
While the additional image is necessary, there is no strict requirement on its degree of difference from the original. Hallucinations typically occur in content-rich elements such as objects, attributes, and relationships, which naturally exhibit substantial variability across real-world images. Furthermore, our experiments demonstrate that the effectiveness of hallucination mitigation is not strongly correlated with image similarity. The experimental results demonstrate that randomly selected images perform comparably to retrieved ones, highlighting the practical applicability of our method. This flexibility significantly lowers the barrier for real-world deployment.
Twice Forward Propagation
Although performing twice forward propagations doubles the theoretical computational cost, in practice, we merge the two inputs into a single batch and perform one forward propagation with a doubled batch size to achieve the same effect. Under this implementation, our method incurs only a 10% increase in inference time and a 5% increase in memory usage, demonstrating strong potential for practical deployment.
Appendix A Detailed Related Work
A.1 Large Vision-Language Models
Typically, Large Vision-Language Models (LVLMs) Dai et al. (2023); Liu et al. (2024c); Bai et al. (2023b); Zhu et al. (2024a) consist of a visual encoder Radford et al. (2021); Dosovitskiy et al. (2021) and a Large Language Model (LLM) backbone Touvron et al. (2023); Chiang et al. (2023). The visual encoder extracts visual features, which are then aligned with text tokens through a fusion module to form the input context for the LLM. Through multi-stage training, including modality alignment and instruction tuning Liu et al. (2023), LVLMs acquire the ability to perform various multimodal tasks such as image captioning Ye et al. (2025), visual question answering Yu et al. (2024), and visual grounding Lu et al. (2024). Despite significant advancements, LVLMs still suffer from severe hallucination problems, often generating content that contradicts the visual input.
A.2 Hallucination in LVLMs
Originally, hallucinations Ji et al. (2023); Liu et al. (2024b) were defined as undesirable outputs or non-factual information. In the multimodal domain, the study of hallucinations primarily focuses on cross-modal consistency Li et al. (2023c); Wang et al. (2023), requiring that LVLM-generated content remains faithful to the visual inputs. Multiple factors contribute to hallucinations in LVLMs, including biased visual attention Zhu et al. (2025); Li et al. (2025), attention misallocation Huang et al. (2024); Yin et al. (2025); Kang et al. (2025), and language priors Leng et al. (2024); Zhu et al. (2024b); Wang et al. (2024c). Among these, language priors are the most direct cause, making it difficult for LVLMs to accurately express their understanding of visual information.
A.3 Language Priors
As LVLMs are built upon LLMs, they inherently inherit the linguistic knowledge embedded in LLMs, which are referred to as language priors Zhibo et al. (2023); Wu et al. (2022). Language priors specifically manifest when LVLMs generate content that is linguistically plausible but inconsistent with visual evidence. Some works Chen et al. (2025); Liu et al. (2024a); Gunjal et al. (2024) meticulously design specialized training data to mitigate this issue. However, such training-based approaches exhibit limited generalizability and cannot be easily transferred to broader application scenarios. Recently, a series of training-free approaches have directly reduced language priors using contrastive decoding Li et al. (2023a); Chuang et al. (2024), achieving remarkable performance. These methods construct an alternative logit distribution on top of the original one through techniques such as masking the image Leng et al. (2024), perturbing the instruction Wang et al. (2024c), augmenting the vision input Zhu et al. (2024b), or performing cross-modal conversion Zhang et al. (2025). During decoding, the two logit distributions are contrasted to eliminate language priors.
Appendix B Detailed Experiment Settings
B.1 Benchmarks
In order to validate the effectiveness of our method in mitigating language hallucination, we evaluate its performance on four widely-used multimodal hallucination benchmarks: two generative benchmarks (CHAIR Rohrbach et al. (2018) and DetailCaps Ye et al. (2025)), one discriminative benchmark (POPE Li et al. (2023b)), and one hybrid benchmark (AMBER Wang et al. (2024a)).
CHAIR
assesses the proportion of hallucinated objects, which are generated by the model but not present in the reference annotations. Following prior works, we randomly select 500 images from the MSCOCO Lin et al. (2014) dataset as the test set. We additionally select another 500 images to construct a validation set for hyperparameter tuning.
DetailCaps
is a fine-grained image captioning benchmark, accompanied by ground-truth detail captions generated by GPT-4V Yang et al. (2023), Gemini-1.5-Pro Team et al. (2024), and GPT-4o OpenAI et al. (2024) for evaluation. It contains 4,870 images from multiple datasets, and we use a subset of 700 images from MSCOCO to conduct our experiments.
POPE
is a widely adopted benchmark for evaluating object hallucinations by prompting LVLMs to identify whether a specific object is present in the image. POPE incorporates three distinct datasets: MSCOCO, A-OKVQA, and GQA. Each dataset uses three different negative sampling settings: Random, Popular, and Adversarial. Finally, it consists of 9,000 questions and 1,500 images.
AMBER
combines generative and discriminative tasks, and is evaluated on a curated set of 1,004 images. In addition to image captioning, it includes 14,216 questions designed to assess hallucinations in object, attribute, and relation recognition.
B.2 Evaluated LVLMs
To demonstrate the generalizability of the proposed CICD as a broadly applicable plug-and-play module, we evaluate its performance across six different LVLMs spanning three model families: InstructBLIP Dai et al. (2023), LLaVA-1.5 Liu et al. (2024c), LLaVA-Next Liu et al. (2024d), Qwen-VL-Chat Bai et al. (2023b), Qwen2-VL-Instruct Wang et al. (2024b), and Qwen2.5-VL-Instruct Bai et al. (2025). InstructBLIP and LLaVA-1.5 are both built upon Vicuna-7B Chiang et al. (2023) as their language backbone, while LLaVA-Next is based on LLaMA3. The Qwen-VL series, including Qwen2-VL-Instruct and Qwen2.5-VL-Instruct, is developed using the QwenLM framework. As for vision encoders, InstructBLIP, LLaVA-1.5, LLaVA-Next, and Qwen-VL-Chat utilize the pretrained CLIP model Radford et al. (2021), whereas Qwen2-VL-Instruct and Qwen2.5-VL-Instruct adopt a fine-tuned ViT Dosovitskiy et al. (2021). All models are implemented at the 8B parameter scale.
B.3 Baselines
We compare our method with several training-free baselines that specifically target language priors.
VCD Leng et al. (2024) introduces noise into the input image to suppress language priors, and then applies contrastive decoding for their elimination.
ICD Wang et al. (2024c) constructs negative logits using adversarial instructions and enhances positive signals through contrastive decoding.
IBD Zhu et al. (2024b) strengthens the model’s attention to visual features, thereby increasing the influence of image information during decoding.
DeFG Zhang et al. (2025) converts the generated caption into an image and re-generates the caption from that image, leveraging the preserved visual content to reduce hallucinations.
Appendix C More Experiment Results
C.1 Effectiveness on Larger LVLMs
We evaluate the effectiveness of our method on larger models using two benchmarks, POPE and CHAIR, with the results presented in Tab.6. Our method continues to deliver significant performance improvements on a 13B model, indicating its strong generalizability. These results suggest that our approach scales well and remains effective across different model sizes.
Model Method CHAIR(MaxLen=512) POPE (F1$\uparrow$)
[…]
LLaVA-1.5-13B Regular 58.8 17 73.4 82.94 81.46 78.49
CICD 44.4 12.9 76.0 87.92 85.68 80.99
INstructBLIP-13B Regular 65.6 21.0 62.7 81.31 77.20 74.69
CICD 44 11.6 65.7 87.52 83.60 78.96
Table 6: Results on larger LVLMs. The maximum generation length is 512.
Method LLaVA-Next Qwen2-VL-Instruct Qwen2.5-VL-Instruct
[…]
Regular 61.06 38.0 10.4 60.04 51.3 10.6 62.62 42.2 11.7
CICD 62.19 36.2 8.5 61.79 48.4 9.9 65.07 43.8 10.9
Table 7: Results on DetailCaps with more LVLMs.